The age of a man is four times that of his son. Five years ago, the man was nine times as old as his son at that time. The present age of the man is

Grade:12

2 Answers

bharat bajaj IIT Delhi
9 years ago
Let the age of son be x.
Hence, theageof father = 4x
Age of son 5 years ago = x - 5
Age of father 5 years ago = 4x - 5
Its given that, 5 years ago the man was 9 times old as his son.
? 4x - 5 = 9 (x - 5)
5x = 40
x = 8
Presentage of son = x = 8 years.
Presentage of father = 4x = 4 × 8 = 32 years.
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty
bharat bajaj IIT Delhi
9 years ago
Let theage of son be x.
Hence, theageof father = 4x
Age of son 5 years ago = x - 5
Age of father 5 years ago = 4x - 5
Its given that, 5 years ago the man was 9 times old as his son.
? 4x - 5 = 9 (x - 5)
5x = 40
x = 8
Presentage of son = x = 8 years.
Presentage of father = 4x = 4 × 8 = 32 years.
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty